# MEM1: Learning to Synergize Memory and Reasoning for Efficient Long-Horizon Agents

### v1

@@ Abstract
Modern language agents must operate over long-horizon, multi-turn interactions, where they retrieve external information, adapt to observations, and answer interdependent queries. Yet, most LLM systems rely on full-context prompting, appending all past turns regardless of their relevance. This leads to unbounded memory growth, increased computational costs, and degraded reasoning performance on out-of-distribution input lengths. We introduce MEM1, an end-to-end reinforcement learning framework that enables agents to operate with constant memory across long multi-turn tasks. At each turn, MEM1 updates a compact shared internal state that jointly supports memory consolidation and reasoning. This state integrates prior memory with new observations from the environment while strategically discarding irrelevant or redundant information. To support training in more realistic and compositional settings, we propose a simple yet effective and scalable approach to constructing multi-turn environments by composing existing datasets into arbitrarily complex task sequences. Experiments across three domains, including internal retrieval QA, open-domain web QA, and multi-turn web shopping, show that MEM1-7B improves performance by while reducing memory usage by compared to Qwen2.5-14B-Instruct on a -objective multi-hop QA task, and generalizes beyond the training horizon. Our results demonstrate the promise of reasoning-driven memory consolidation as a scalable alternative to existing solutions for training long-horizon interactive agents, where both efficiency and performance are optimized.
@@
As illustrated in Fig.˜1, at each turn, the model updates a consolidated state composed of prior memory and newly obtained information. This consolidated state becomes the agent’s only retained memory, allowing all external tool outputs to be discarded after use, which prevents prompt expansion altogether as illustrated by Fig.˜2. A key insight of our method is that inference-time reasoning [52, 13, 30, 62] serves a dual function: it not only provides deeper insight into the current query but also acts as a form of “working memory” [4], extracting key components from gathered information to build an evolving understanding. By unifying reasoning and memory consolidation, MEM1 offers an elegant solution where the agent learns to reason and remember within a shared representational space without requiring additional modules or architectural changes.
@@ 3.1 Memory as Part of Reasoning
To achieve a constant memory, MEM1 is particularly trained to iteratively refine its understanding by processing new information in conjunction with a consolidation of its prior state. At each turn , the agent produces a new <IS_t> element, which summarizes past information and reasons about subsequent actions. Following this, the agent generates an action—either a <query_t> to interact with the environment, or an <answer_t> if a direct response is warranted. If the agent issues a <query_t>, the corresponding feedback from the environment is appended as <info_t>. At the next turn, , the agent consolidates the tuple into a new <IS_(t+1)>, which serves as the basis for further interactions. After each turn, all tags from the previous turn are pruned from the context, effectively compressing memory and preventing prompt bloat. Fig.˜3 (bottom left) illustrates the evolution of the model’s context over time. At any given turn, the agent retains at most two <IS> elements, two <query> elements, and one <info> element, ensuring bounded and efficient memory usage. The detailed rollout algorithm is in Alg.˜1 of Sec.˜A.5.
RL offers a powerful mechanism for shaping agent behavior through reward signals [46]. In MEM1, we leverage this framework to incentivize effective state consolidation by designing environments in which the agent is rewarded only when it strategically retains and integrates useful information. Specifically, we construct tasks that require numerous interactions with the environment to arrive at a correct answer (see Sec.˜3.3). Success depends on the agent’s ability to rely on information collected along the inference path. At each turn, we prune the agent’s context to retain only the most recent <IS>, forcing the agent to perform memory consolidation as part of its reasoning process. Without access to full historical context, the agent must learn to preserve and update relevant knowledge internally in order to reap the reward. This learning procedure mirrors how humans cultivate memorization skills through structured tasks such as Sudoku or crosswords [1], where success hinges on selectively attending to key information and building upon it. Over time, such tasks help individuals develop cognitive strategies that jointly support efficient memorization and reasoning, similar to our RL method for training MEM1.
@@ 3.2 Masked Trajectory for Policy Optimization
To address this, we introduce a masked trajectory approach that reconstructs a logically coherent full trajectory by stitching together multiple interaction turns with evolving contexts. This unified trajectory mimics a standard multi-turn rollout and comprises a sequence of tuples for , where denotes the total number of interaction turns. The -th turn outputs the final answer . The full trajectory encodes all information needed for accurate policy learning while respecting MEM1’s memory consolidation at each turn. Fig.˜3 (bottom left) demonstrates the evolution of the agent’s context.
To ensure that policy gradients are correctly computed under this consolidated memory regime, we apply a two-dimensional attention mask [37] across the full trajectory. This mask restricts each token’s attention to only the tokens retained in memory at the time that token was generated. Specifically, for a token position , we mask out all prior tokens that are not part of the consolidated memory corresponding to the context at that turn. This masking strategy enables accurate computation of the policy objective: letting denote the masked input state and the predicted token, the log-probability ratio remains valid and tractable, in turn ensuring that the advantage, KL penalty, and value estimation are correct. A visualization is in Fig.˜3 (bottom right). Furthermore, following [22], we incorporate a one-dimensional attention mask over retrieved external information during policy updates for both the actor and critic networks. This ensures that gradient updates are localized to only tokens generated by the agent. Fig.˜3 (bottom right) shows the masking mechanism that enables stable and accurate policy optimization under MEM1’s memory-constrained execution.
@@ 3.3 Multi-Objective Task Design
To bridge this gap, we introduce a novel task—multi-objective question & answering (QA)—that extends the number of reasoning steps required to solve a problem. Building on existing multi-turn datasets such as HotpotQA and Natural Question [56, 24], we interleave multiple questions from the original QA corpus and construct a single composite query that requires answering all constituent sub-questions, shown in Prompt A.3 of Sec.˜A.3. This formulation compels the agent to perform multiple search queries, each targeting a distinct sub-question or sub-objective, and then integrate the retrieved answers to form a comprehensive final response. Compared to the original tasks, our synthesized multi-objective, multi-hop setting significantly increases the number of search and reasoning turns required, leading to more complex, memory-intensive interactions.
@@ 4 Experiments & Results
Our experiments are conducted in two standard environments, each reflecting real-world scenarios that require multi-turn agent interactions. The first environment is question answering with retrieval-augmented generation (RAG) [24, 56], where the agent must answer queries by retrieving relevant information from an external knowledge store (either a database or an online search engine). We trained on RAG with a local database (i.e., Wikipedia Corpus) and evaluated on tasks involving open web browsing. For QA, following Sec.˜3.3, we construct multi-objective tasks and tested the model performance on tasks with more questions than seen in the training. The second environment is WebShop navigation [57], where the agent assists users in online shopping by browsing a website and selecting items based on natural language descriptions. This task requires the agent to iteratively read page content and make navigation decisions, following protocols similar to those in WebGPT [32].
@@ 4.1 Implementation Details
For the web agent, we use the WebShop environment [57], which also produces a reward during training [61]. For all datasets, the train-test split follows the original papers. During RL training, we employ the exact match (EM) metric for QA tasks (details in Sec.˜A.4.1) and the environment reward for WebShop [57, 61]. To evaluate the effectiveness of various approaches, we measure the EM and F1 score for QA tasks and final reward for the WebShop environment [57, 61]. To evaluate the efficiency, we consider the peak token usage, average dependency, and average inference time. The test datasets are obtained from the original papers which consist of out-of-distribution data. The former two metrics measure the memory efficiency, while the latter measures the time efficiency. The detailed definitions of the metrics are in Sec.˜A.4.1. The prompt and format can be found in Sec.˜A.3.
@@
To evaluate the accuracy and efficiency of MEM1, we compare it against a diverse set of baselines designed to either enhance task performance or manage context effectively. For the QA environment, we benchmark accuracy against Search-R1 [22], DeepResearcher [66], and a larger-scale model, Qwen2.5-14B-Instruct [55]. Details about Search-R1 and DeepResearcher can be found in Sec.˜A.4.2. For the WebShop environment, we compare against Agent-FLAN [9], Agent-R [61], and AgentLM [63]. To assess efficiency, we consider two context compression baselines using models of the same parameter size as MEM1. First, we apply MEM1’s agentic truncation prompt template and rollout to a standard instruct model, isolating the benefits of prompt and rollout design alone. Second, we evaluate A-MEM [54], which augments an Instruct model with a vector database for memory retrieval, capturing the effect of external memory modules in agentic systems. We additionally train a supervised fine-tuned (SFT) model using trajectories curated from GPT-4o [33] based on MEM1’s rollout and compare it with the RL-trained agent.
@@
In our agentic pipeline, the agent’s context is programmatically truncated at each turn—immediately after it generates a search query or an answer—following the procedure outlined in Sec.˜3. As past context is truncated, the agent may have difficulty determining when to terminate. To address this, we prepend a hint [HINT: YOU HAVE {turns_left} TURNS LEFT] at the beginning of each <info> tag to remind the agent of its remaining turns budget. For all experiments, we set the maximally allowed turns to for -objective to -objective tasks and for more difficult tasks to avoid excessively long trajectories.
@@ 4.2 MEM1 on Multi-Objective Multi-Hop Tasks
One key advantage of MEM1 agents lies in their efficient management of long-horizon interactions with the environment. To demonstrate this, we train our MEM1 agent with a 2-objective augmentation of the QA dataset, and subsequently test it against other models, using held-out multi-objective test datasets similarly augmented from the original test datasets. As elaborated in Sec.˜3.3, these multi-objective tasks require a significantly larger number of turns of environment interactions to complete, hence serving as better benchmarks for memory management. As shown in Tab.˜1, when evaluated on 2-objective datasets, MEM1 achieves better performance (in terms of EM and F1 scores) than other 7B counterparts, while incurring significantly lower peak token usage and achieving faster inference time.
The advantage of MEM1 becomes even more evident in tasks requiring longer-horizon interactive processes. To highlight such scalability of MEM1, we further compare the models on , and -objective tasks in Fig.˜4 and Tab.˜1. Fig.˜4 illustrates the scaling trends of task performance (measured by EM count) and memory efficiency (measured by Peak Token Usage) for MEM1 relative to other models and memory management baselines. As the number of objectives increases, the Peak Token Usage of all other methods and models scales nearly linearly. In contrast, MEM1 maintains an almost constant peak token count with only a slight increase, as also shown in Tab.˜1.
@@ 4.3 MEM1 on Single-Objective Multi-Hop Tasks
Beyond QA tasks, we further evaluate the effectiveness of MEM1 in managing long-horizon interactions in the form of web navigation. We show the experimental results in Tab.˜2. Trained in the WebShop environment (see Sec.˜A.6), MEM1 outperforms other agent training baselines, including Agent-Flan, Agent-R, and AgentLM when utilizing models of similar size. Furthermore, MEM1 achieves remarkable efficiency improvements compared to the best baseline method, AgentLM, featuring a improvement in Peak Token Usage, a improvement in Dependency, and a improvement in Inference Time. MEM1 even surpasses AgentLM-13B, a model with twice the parameter count of our trained model. Additionally, our results indicate that using MEM1 is significantly better than OpenAI’s GPT-4o on the WebShop tasks, even when the truncation prompt templates or A-MEM techniques are applied to GPT-4o.
@@
Tab.˜3 presents the accuracy and efficiency metrics for evaluations on single-objective QA tasks on Wikipedia [22], where the agent can make retrieval requests from the Wikipedia datastore via RAG. The MEM1 used in this evaluation is the same as the one detailed in Sec.˜4.2, which is trained solely on a 2-objective task. Overall, MEM1 demonstrates superior efficiency across all three evaluated efficiency metrics, while simultaneously achieving the highest EM score and an F1 score comparable to that of Qwen2.5-14B-Instruct. This improvement in efficiency is attributed to the MEM1 agent’s ability to consolidate memory from previous interactions into a compact internal state, which reduces the number of tokens used in the context. We also observe that SFT significantly underperforms RL, highlighting the necessity for RL-based training.
@@
To validate the transferability and generalizability of the trained MEM1 agent, we perform a zero-shot transfer to an online web-QA environment, which is unseen by the agent. In this environment, agents conduct web searches through an API service that returns results including titles, snippets, and URLs. As shown in Tab.˜3, MEM1 consistently exhibited improved efficiency alongside comparable effectiveness in this unseen setting via zero-shot transfer.
@@ 4.4 Analysis on Emergent Agent Behaviors
Through analyzing MEM1’s multi-turn interaction traces trained on 2-objective QA, we observe a range of emergent behaviors that are critical for handling long-horizon, multi-objective tasks, demonstrating capabilities well beyond simple retrieval. First, MEM1 learns to manage multiple questions concurrently by maintaining a structured internal state. As shown in Fig.˜5(a), when faced with two multi-hop questions, the agent stores and updates memory for each question separately, guiding subsequent searches based on the identified information gaps. In (b), MEM1 exhibits the ability to shift focus when progress on one question stalls, recognizing difficulty and prioritizing the more tractable objective. Meanwhile, MEM1 learns to interleave reasoning and memory in its internal state <IS>, weaving important information into its decision-making process to support both information retention and action selection. In Fig.˜5 (c), MEM1 explicitly extracts important information from previous search results and leverages it to formulate the next query that best addresses the current information gap. In addition, (d) shows that when new, relevant information is retrieved, MEM1 explicitly reasons about its significance and selectively updates its memory. We believe that learning these interleaved behaviors is key to achieving efficiency gains in memory without degrading performance. Beyond behaviors unique to our multi-objective setup and memory architecture, MEM1 also exhibits several general-purpose search strategies. In (e), the agent performs self-verification, correcting an earlier misconception and issuing a new query for confirmation. In (f), complex queries are decomposed into manageable subgoals before initiating the search. In (g), for questions requiring multi-turn information gathering, MEM1 extracts key information from search results and uses it to inform the next search. In (h), when overly specific queries fail, MEM1 re-scopes its query to improve retrieval. Notably, many of these behaviors, including verification, making a plan, and iterative search, are also reported in recent studies on deep research agents [22, 66].

### v2

@@ Abstract
Modern language agents must operate over long-horizon, multi-turn interactions, where they retrieve external information, adapt to observations, and answer interdependent queries. Yet, most LLM systems rely on full-context prompting, appending all past turns regardless of their relevance. This leads to unbounded memory growth, increased computational costs, and degraded reasoning performance on out-of-distribution input lengths. We introduce MEM1, an end-to-end reinforcement learning framework that enables agents to operate with constant memory across long multi-turn tasks. At each turn, MEM1 updates a compact shared internal state that jointly supports memory consolidation and reasoning. This state integrates prior memory with new observations from the environment while strategically discarding irrelevant or redundant information. To support training in more realistic and compositional settings, we propose a simple yet effective and scalable approach to constructing multi-turn environments by composing existing datasets into arbitrarily complex task sequences. Experiments across three domains, including internal retrieval QA, open-domain web QA, and multi-turn web shopping, show that MEM1-7B improves performance by while reducing memory usage by compared to Qwen2.5-14B-Instruct on a -objective multi-hop QA task, and generalizes beyond the training horizon. Our results demonstrate the promise of reasoning-driven memory consolidation as a scalable alternative to existing solutions for training long-horizon interactive agents, where both efficiency and performance are optimized. Code can be found at https://github.com/MIT-MI/MEM1.
@@
As illustrated in Fig. 1, at each turn, the model updates a consolidated state composed of prior memory and newly obtained information. This consolidated state becomes the agent’s only retained memory, allowing all external tool outputs to be discarded after use, which prevents prompt expansion altogether as illustrated by Fig. 2. A key insight of our method is that inference-time reasoning [55, 15, 33, 65] serves a dual function: it not only provides deeper insight into the current query but also acts as a form of “working memory” [4], extracting key components from gathered information to build an evolving understanding. By unifying reasoning and memory consolidation, MEM1 offers an elegant solution where the agent learns to reason and remember within a shared representational space without requiring additional modules or architectural changes.
@@ 3.1 Memory as Part of Reasoning
To achieve a constant memory, MEM1 is particularly trained to iteratively refine its understanding by processing new information in conjunction with a consolidation of its prior state. At each turn , the agent produces a new <IS_t> element, which summarizes past information and reasons about subsequent actions. Following this, the agent generates an action—either a <query_t> to interact with the environment, or an <answer_t> if a direct response is warranted. If the agent issues a <query_t>, the corresponding feedback from the environment is appended as <info_t>. At the next turn, , the agent consolidates the tuple into a new <IS_(t+1)>, which serves as the basis for further interactions. After each turn, all tags from the previous turn are pruned from the context, effectively compressing memory and preventing prompt bloat. Fig. 3 (bottom left) illustrates the evolution of the model’s context over time. At any given turn, the agent retains at most two <IS> elements, two <query> elements, and one <info> element, ensuring bounded and efficient memory usage. The detailed rollout algorithm is in Alg. 1 of Sec. A.5.
RL offers a powerful mechanism for shaping agent behavior through reward signals [49]. In MEM1, we leverage this framework to incentivize effective state consolidation by designing environments in which the agent is rewarded only when it strategically retains and integrates useful information. Specifically, we construct tasks that require numerous interactions with the environment to arrive at a correct answer (see Sec. 3.3). Success depends on the agent’s ability to rely on information collected along the inference path. At each turn, we prune the agent’s context to retain only the most recent <IS>, forcing the agent to perform memory consolidation as part of its reasoning process. Without access to full historical context, the agent must learn to preserve and update relevant knowledge internally in order to reap the reward. This learning procedure mirrors how humans cultivate memorization skills through structured tasks such as Sudoku or crosswords [1], where success hinges on selectively attending to key information and building upon it. Over time, such tasks help individuals develop cognitive strategies that jointly support efficient memorization and reasoning, similar to our RL method for training MEM1.
@@ 3.2 Masked Trajectory for Policy Optimization
To address this, we introduce a masked trajectory approach that reconstructs a logically coherent full trajectory by stitching together multiple interaction turns with evolving contexts. This unified trajectory mimics a standard multi-turn rollout and comprises a sequence of tuples for , where denotes the total number of interaction turns. The -th turn outputs the final answer . The full trajectory encodes all information needed for accurate policy learning while respecting MEM1’s memory consolidation at each turn. Fig. 3 (bottom left) demonstrates the evolution of the agent’s context.
To ensure that policy gradients are correctly computed under this consolidated memory regime, we apply a two-dimensional attention mask [40] across the full trajectory. This mask restricts each token’s attention to only the tokens retained in memory at the time that token was generated. Specifically, for a token position , we mask out all prior tokens that are not part of the consolidated memory corresponding to the context at that turn. This masking strategy enables accurate computation of the policy objective: letting denote the masked input state and the predicted token, the log-probability ratio remains valid and tractable, in turn ensuring that the advantage, KL penalty, and value estimation are correct. A visualization is in Fig. 3 (bottom right). Furthermore, following [24], we incorporate a one-dimensional attention mask over retrieved external information during policy updates for both the actor and critic networks. This ensures that gradient updates are localized to only tokens generated by the agent. Fig. 3 (bottom right) shows the masking mechanism that enables stable and accurate policy optimization under MEM1’s memory-constrained execution.
@@ 3.3 Multi-Objective Task Design
To bridge this gap, we introduce a novel task—multi-objective question & answering (QA)—that extends the number of reasoning steps required to solve a problem. Building on existing multi-turn datasets such as HotpotQA and Natural Question [59, 26], we interleave multiple questions from the original QA corpus and construct a single composite query that requires answering all constituent sub-questions, shown in Prompt A.3 of Sec. A.3. This formulation compels the agent to perform multiple search queries, each targeting a distinct sub-question or sub-objective, and then integrate the retrieved answers to form a comprehensive final response. Compared to the original tasks, our synthesized multi-objective, multi-hop setting significantly increases the number of search and reasoning turns required, leading to more complex, memory-intensive interactions.
@@ 4 Experiments & Results
Our experiments are conducted in two standard environments, each reflecting real-world scenarios that require multi-turn agent interactions. The first environment is question answering with retrieval-augmented generation (RAG) [26, 59], where the agent must answer queries by retrieving relevant information from an external knowledge store (either a database or an online search engine). We trained on RAG with a local database (i.e., Wikipedia Corpus) and evaluated on tasks involving open web browsing. For QA, following Sec. 3.3, we construct multi-objective tasks and tested the model performance on tasks with more questions than seen in the training. The second environment is WebShop navigation [60], where the agent assists users in online shopping by browsing a website and selecting items based on natural language descriptions. This task requires the agent to iteratively read page content and make navigation decisions, following protocols similar to those in WebGPT [35].
@@ 4.1 Implementation Details
For the web agent, we use the WebShop environment [60], which also produces a reward during training [64]. For all datasets, the train-test split follows the original papers. During RL training, we employ the exact match (EM) metric for QA tasks (details in Sec. A.4.1) and the environment reward for WebShop [60, 64]. To evaluate the effectiveness of various approaches, we measure the EM and F1 score for QA tasks and final reward for the WebShop environment [60, 64]. To evaluate the efficiency, we consider the peak token usage, average dependency, and average inference time. The test datasets are obtained from the original papers which consist of out-of-distribution data. The former two metrics measure the memory efficiency, while the latter measures the time efficiency. The detailed definitions of the metrics are in Sec. A.4.1. The prompt and format can be found in Sec. A.3.
@@
To evaluate the accuracy and efficiency of MEM1, we compare it against a diverse set of baselines designed to either enhance task performance or manage context effectively. For the QA environment, we benchmark accuracy against Search-R1 [24], DeepResearcher [69], and a larger-scale model, Qwen2.5-14B-Instruct [58]. Details about Search-R1 and DeepResearcher can be found in Sec. A.4.2. For the WebShop environment, we compare against Agent-FLAN [11], Agent-R [64], and AgentLM [66]. To assess efficiency, we consider two context compression baselines using models of the same parameter size as MEM1. First, we apply MEM1’s agentic truncation prompt template and rollout to a standard instruct model, isolating the benefits of prompt and rollout design alone. Second, we evaluate A-MEM [57], which augments an Instruct model with a vector database for memory retrieval, capturing the effect of external memory modules in agentic systems. We additionally train a supervised fine-tuned (SFT) model using trajectories curated from GPT-4o [36] based on MEM1’s rollout and compare it with the RL-trained agent.
@@
In our agentic pipeline, the agent’s context is programmatically truncated at each turn—immediately after it generates a search query or an answer—following the procedure outlined in Sec. 3. As past context is truncated, the agent may have difficulty determining when to terminate. To address this, we prepend a hint [HINT: YOU HAVE {turns_left} TURNS LEFT] at the beginning of each <info> tag to remind the agent of its remaining turns budget. For all experiments, we set the maximally allowed turns to for -objective to -objective tasks and for more difficult tasks to avoid excessively long trajectories.
@@ 4.2 MEM1 on Multi-Objective Multi-Hop Tasks
One key advantage of MEM1 agents lies in their efficient management of long-horizon interactions with the environment. To demonstrate this, we train our MEM1 agent with a 2-objective augmentation of the QA dataset, and subsequently test it against other models, using held-out multi-objective test datasets similarly augmented from the original test datasets. As elaborated in Sec. 3.3, these multi-objective tasks require a significantly larger number of turns of environment interactions to complete, hence serving as better benchmarks for memory management. As shown in Tab. 1, when evaluated on 2-objective datasets, MEM1 achieves better performance (in terms of EM and F1 scores) than other 7B counterparts, while incurring significantly lower peak token usage and achieving faster inference time.
The advantage of MEM1 becomes even more evident in tasks requiring longer-horizon interactive processes. To highlight such scalability of MEM1, we further compare the models on , and -objective tasks in Fig. 4 and Tab. 1. Fig. 4 illustrates the scaling trends of task performance (measured by EM count) and memory efficiency (measured by Peak Token Usage) for MEM1 relative to other models and memory management baselines. As the number of objectives increases, the Peak Token Usage of all other methods and models scales nearly linearly. In contrast, MEM1 maintains an almost constant peak token count with only a slight increase, as also shown in Tab. 1.
@@ 4.3 MEM1 on Single-Objective Multi-Hop Tasks
Beyond QA tasks, we further evaluate the effectiveness of MEM1 in managing long-horizon interactions in the form of web navigation. We show the experimental results in Tab. 2. Trained in the WebShop environment (see Sec. A.6), MEM1 outperforms other agent training baselines, including Agent-Flan, Agent-R, and AgentLM when utilizing models of similar size. Furthermore, MEM1 achieves remarkable efficiency improvements compared to the best baseline method, AgentLM, featuring a improvement in Peak Token Usage, a improvement in Dependency, and a improvement in Inference Time. MEM1 even surpasses AgentLM-13B, a model with twice the parameter count of our trained model. Additionally, our results indicate that using MEM1 is significantly better than OpenAI’s GPT-4o on the WebShop tasks, even when the truncation prompt templates or A-MEM techniques are applied to GPT-4o.
@@
Tab. 3 presents the accuracy and efficiency metrics for evaluations on single-objective QA tasks on Wikipedia [24], where the agent can make retrieval requests from the Wikipedia datastore via RAG. The MEM1 used in this evaluation is the same as the one detailed in Sec. 4.2, which is trained solely on a 2-objective task. Overall, MEM1 demonstrates superior efficiency across all three evaluated efficiency metrics, while simultaneously achieving the highest EM score and an F1 score comparable to that of Qwen2.5-14B-Instruct. This improvement in efficiency is attributed to the MEM1 agent’s ability to consolidate memory from previous interactions into a compact internal state, which reduces the number of tokens used in the context. We also observe that SFT significantly underperforms RL, highlighting the necessity for RL-based training.
@@
To validate the transferability and generalizability of the trained MEM1 agent, we perform a zero-shot transfer to an online web-QA environment, which is unseen by the agent. In this environment, agents conduct web searches through an API service that returns results including titles, snippets, and URLs. As shown in Tab. 3, MEM1 consistently exhibited improved efficiency alongside comparable effectiveness in this unseen setting via zero-shot transfer.
@@ 4.4 Analysis on Emergent Agent Behaviors
Through analyzing MEM1’s multi-turn interaction traces trained on 2-objective QA, we observe a range of emergent behaviors that are critical for handling long-horizon, multi-objective tasks, demonstrating capabilities well beyond simple retrieval. First, MEM1 learns to manage multiple questions concurrently by maintaining a structured internal state. As shown in Fig. 5(a), when faced with two multi-hop questions, the agent stores and updates memory for each question separately, guiding subsequent searches based on the identified information gaps. In (b), MEM1 exhibits the ability to shift focus when progress on one question stalls, recognizing difficulty and prioritizing the more tractable objective. Meanwhile, MEM1 learns to interleave reasoning and memory in its internal state <IS>, weaving important information into its decision-making process to support both information retention and action selection. In Fig. 5 (c), MEM1 explicitly extracts important information from previous search results and leverages it to formulate the next query that best addresses the current information gap. In addition, (d) shows that when new, relevant information is retrieved, MEM1 explicitly reasons about its significance and selectively updates its memory. We believe that learning these interleaved behaviors is key to achieving efficiency gains in memory without degrading performance. Beyond behaviors unique to our multi-objective setup and memory architecture, MEM1 also exhibits several general-purpose search strategies. In (e), the agent performs self-verification, correcting an earlier misconception and issuing a new query for confirmation. In (f), complex queries are decomposed into manageable subgoals before initiating the search. In (g), for questions requiring multi-turn information gathering, MEM1 extracts key information from search results and uses it to inform the next search. In (h), when overly specific queries fail, MEM1 re-scopes its query to improve retrieval. Notably, many of these behaviors, including verification, making a plan, and iterative search, are also reported in recent studies on deep research agents [24, 69].